Gucchi:

Simplify negative 8 over 2 ÷ 6 over negative 3. -4 2 −2 4

Gucchi:

im thinking its 2

Gucchi:

just wanna make sure with y'all

umm:

\[\frac{ -8 }{ 2}\div \frac{ 6 }{ -3 }\]

Gucchi:

yh it has to be 2

supie:

ohh lemme write that again then.

supie:

\[8/-2÷6/-3=\] \[-4/÷(-2)=2\]